# Debiased Inference for Dynamic Nonlinear Panels with Multi-dimensional Heterogeneities

### v3

@@ 3 Asymptotic Theory
    $a_{i}\left(m\right):=\sup_{\theta\in\Theta}\sup_{t\geq 1}\sup_{A\in\mathcal{A}_{it},B\in\mathcal{B}_{it+m}}\left|\mathbb{P}\left(A\cap B\mid\phi(\theta)\right)-\mathbb{P}\left(A\mid\phi(\theta)\right)\mathbb{P}\left(B\mid\phi(\theta)\right)\right|,$
    and, $\sup_{N}\max_{1\leq i\leq N}\sum_{m=0}^{\infty}\left[a_{i}\left(m\right)\right]^{1-2/\eta}<\infty,$ for some $\eta>2.$
  • (vi) For $\theta\in\Theta,$ $\sqrt{NT}\,\mathbb{E}\left[{H}(\theta)\right]$ is nonsingular when $N$ and $T$ are large.
Assumption 1 requires that both $N$ and $T$ are large and proportional, which is standardis standard in the literature. Assumption 1 imposes smoothness and boundedness conditions on the likelihood. Specifically, it requires that the likelihood and its derivatives for unit $i$ at time $t$ are unformly bounded by a measurable function whose moments of order greater than 2 exist. Assumption 1 is the identification condition for the incidental parameters (fixed effects) as they enter the likelihood additively. Given Assumption 1, Assumption 1 ensures that for each fixed $\theta,$ $\psi(\theta)$ defined in (2.2) is the unique maximizer. Assumption 1 assumes that the data are independent across units $i$ and conditional strong mixing across time $t.$ Under Assumptions 1 and 1, we can apply Davydov’s inequality to control the convergence rate of certain expectations of summations. Assumption 1 requires that the identified incidental parameter Hessian is invertible.
The following lemma states the bias induced by IPP, as shown in (2.7).
Lemma 1
Under Assumption 1, we have
$\mathbb{E}\left\{s^{\prime}(\theta)[{H}\left(\theta\right)]^{-1}s(\theta)\right\}=\mathbb{E}\left\{s_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)s_{\breve{\alpha}}(\theta)\right\}+\mathbb{E}\left\{s_{\breve{\gamma}}^{\prime}(\theta)H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)s_{\breve{\gamma}}(\theta)\right\}+o\big{(}1/\sqrt{NT}\big{)},$
where $o(\cdot)$ is uniform in $\theta,$ $s_{\breve{\alpha}}(\theta)$ represent the first $(N-1)K$ elements of $s(\theta),$ and $s_{\breve{\gamma}}(\theta)$ represent the last $(T-1)K$ elements of $s(\theta).$
The following theorem presents the bias terms of $\mathbb{E}[\widehat{l}(\theta)]$ to approximate $\mathbb{E}[l(\theta)].$
@@
Under Assumption 1, as $N$ and $T\rightarrow\infty,$ we have
$\mathbb{E}[l(\theta)]=\mathbb{E}[\widehat{l}(\theta)]+B_{\breve{\alpha}}(\theta)+B_{\breve{\gamma}}(\theta)+o\big{(}1/\sqrt{NT}\big{)},$
where $o(\cdot)$ is uniform over $\theta\in\Theta.$ The bias terms $B_{\breve{\alpha}}(\theta)$ and $B_{\breve{\gamma}}(\theta)$ are defined as follows:
@@
$B_{\breve{\alpha}}(\theta):=\frac{1}{2}\mathbb{E}\big{[}{s}_{\breve{\alpha}}^{\prime}(\theta)H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta){s}_{\breve{\alpha}}(\theta)\big{]},\quad B_{\breve{\gamma}}(\theta):=\frac{1}{2}\mathbb{E}\big{[}{s}_{\breve{\gamma}}^{\prime}(\theta){H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta){s}_{\breve{\gamma}}(\theta)\big{]}.$
Since $B_{\breve{\alpha}}(\theta)=O\big{(}1/\sqrt{NT}\big{)}$ and $B_{\breve{\gamma}}(\theta)=O\big{(}1/\sqrt{NT}\big{)}$ as $N$ and $T\to\infty,$ these bias terms introduce a non-zero limiting center to the MLE based on $\widehat{l}(\theta).$ Therefore, it is necessary to eliminate the effect of this bias. In the following, we estimate these bias terms and add them back to $\widehat{l}(\theta).$ Let $\psi(\theta)=:(\alpha^{\prime}_{1}(\theta),\ldots,\alpha^{\prime}_{N-1}(\theta),\gamma^{\prime}_{1}(\theta),\ldots,\gamma^{\prime}_{T-1}(\theta))^{\prime},$ $-\sum_{i=1}^{N-1}\alpha_{i}(\theta)=:\alpha_{N}(\theta),$ and $-\sum_{t=1}^{T-1}\gamma_{t}(\theta)=:\gamma_{T}(\theta).$ Then, $\phi(\theta)=(\alpha^{\prime}_{1}(\theta),\ldots,\alpha^{\prime}_{N}(\theta),\gamma^{\prime}_{1}(\theta),\ldots,\gamma^{\prime}_{T}(\theta))^{\prime}.$ Noting that $\mathbb{E}[s_{\breve{\alpha}}(\theta)]=0$ and $\mathbb{E}[s_{\breve{\gamma}}(\theta)]=0,$ we have
$B_{\breve{\alpha}}(\theta)$ $=\frac{1}{2}\mathbb{E}\big{\{}[s_{\breve{\alpha}}(\theta)-\mathbb{E}s_{\breve{\alpha}}(\theta)]^{\prime}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)[s_{\breve{\alpha}}(\theta)-\mathbb{E}s_{\breve{\alpha}}(\theta)]\big{\}}$
$=\frac{1}{2}\text{{trace}}\big{\{}H_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\alpha}}(\theta)\big{\}},$ (3.1)
where
$\Sigma^{\breve{\alpha}}(\theta)$ $:=\Sigma_{1}^{\breve{\alpha}}(\theta)+\Sigma_{2}^{\breve{\alpha}}(\theta),$
$\Sigma_{1}^{\breve{\alpha}}(\theta)$ $:=\text{{diag}}\big{\{}\mathbb{E}\big{[}\widetilde{s}_{1}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{1}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]},\ldots,\mathbb{E}\big{[}\widetilde{s}_{N-1}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{N-1}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]}\big{\}},$
$\Sigma_{2}^{\breve{\alpha}}(\theta)$ $:=(\iota_{N-1}\iota_{N-1}^{\prime})\otimes\mathbb{E}\big{[}\widetilde{s}_{N}^{\alpha}\left(\theta\right)\left(\widetilde{s}_{N}^{\alpha}\left(\theta\right)\right)^{\prime}\big{]},$
$\widetilde{s}_{i}^{\alpha}\left(\theta\right)$ :=$:=\big{\{}\partial_{\alpha_{i}}l\left(\theta,\phi\right)-\mathbb{E}[\partial_{\alpha_{i}}l\left(\theta,\phi\right)]\big{\}}\big{|}_{\phi=\phi\left(\theta\right)},\quad\text{$i=1,\ldots,N.$}$
Similarly, we have
$B_{\breve{\gamma}}(\theta)$ $=\frac{1}{2}\mathbb{E}\big{\{}[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]^{\prime}H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}\left(\theta\right)[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]\big{\}}$
$=\frac{1}{2}\text{{trace}}\big{\{}H_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\Sigma^{\breve{\gamma}}(\theta)\big{\}},$ (3.2)
where
$\Sigma^{\breve{\gamma}}(\theta)$ $:=\mathbb{E}\big{\{}[s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)][s_{\breve{\gamma}}(\theta)-\mathbb{E}s_{\breve{\gamma}}(\theta)]^{\prime}\big{\}}=\mathbb{E}\big{[}\widetilde{S}_{\breve{\gamma}}(\theta)(\widetilde{S}_{\breve{\gamma}}(\theta))^{\prime}\big{]}$
$\widetilde{S}_{\breve{\gamma}}(\theta)$ $:=\big{(}(\widetilde{s}_{1}^{\gamma}(\theta))^{\prime},\ldots,(\widetilde{s}_{T-1}^{\gamma}(\theta))^{\prime}\big{)}^{\prime}-\iota_{T-1}\otimes\widetilde{s}_{T}^{\gamma}(\theta)$
$\widetilde{s}_{t}^{\gamma}(\theta)$ :=$:=\big{\{}\partial_{\gamma_{t}}l\left(\theta,\phi\right)-\mathbb{E}[\partial_{\gamma_{t}}l\left(\theta,\phi\right)]\big{\}}\big{|}_{\phi=\phi(\theta)}.$
Next, we estimate the two bias terms. We first estimate the Hessian matrix $H(\theta)=H(\theta,\psi(\theta))$ by substituting $\widehat{\psi}(\theta).$ Define $\widehat{H}(\theta):=H(\theta,\widehat{\psi}(\theta)).$ We have $\widehat{H}(\theta)=D\big{[}\partial_{\phi\phi^{\prime}}l(\theta,\phi)\mid_{\phi=\widehat{\phi}(\theta)}\big{]}D^{\prime}$ with $\widehat{\phi}(\theta)=D^{\prime}\widehat{\psi}(\theta).$ Following (3) and (3), we estimate ${B}_{\breve{\alpha}}(\theta)$ and ${B}_{\breve{\gamma}}(\theta)$ by
$\widehat{B}_{\breve{\alpha}}(\theta):=\frac{1}{2}\operatorname*{trace}\left(\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right)\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)\right)\quad\text{and}\quad\widehat{B}_{\breve{\gamma}}(\theta):=\frac{1}{2}\operatorname*{trace}\left(\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\widehat{\Sigma}_{\nu}^{\breve{\gamma}}(\theta)\right),$
where $\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}(\theta)$ and $\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)$ are the first $(N-1)K\times(N-1)K$ and last $(T-1)K\times(T-1)K$ blocks of
@@
$[\widehat{H}(\theta)]^{-1}=:\begin{pmatrix}\widehat{H}_{\breve{\alpha}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right),&\widehat{H}_{\breve{\alpha}\breve{\gamma}^{\prime}}^{\ast}(\theta)\\ \widehat{H}_{\breve{\gamma}\breve{\alpha}^{\prime}}^{\ast}\left(\theta\right),&\widehat{H}_{\breve{\gamma}\breve{\gamma}^{\prime}}^{\ast}(\theta)\end{pmatrix},$
and $\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)$ and $\widehat{\Sigma}_{\nu}^{\breve{\gamma}}(\theta)$ are the estimators of $\Sigma^{\breve{\alpha}}(\theta)$ and $\Sigma^{\breve{\gamma}}(\theta),$ defined as follows. For some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$,
$\widehat{\Sigma}_{\tau}^{\breve{\alpha}}(\theta)$ $:=\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)+\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta),$
@@
$\widehat{\Sigma}_{1,\tau}^{\breve{\alpha}}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}\operatorname*{diag}\big{\{}\widehat{s}_{1,t}^{\alpha}(\theta)\big{(}\widehat{s}_{1,s}^{\alpha}(\theta)\big{)}^{\prime},\ldots,\widehat{s}_{N-1,t}^{\alpha}(\theta)\big{(}\widehat{s}_{N-1,s}^{\alpha}(\theta)\big{)}^{\prime}\big{\}},$
@@
$\widehat{\Sigma}_{2,\tau}^{\breve{\alpha}}(\theta)$ $:=\left(\iota_{N-1}\iota_{N-1}^{\prime}\right)\otimes\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\left\{\left|t-s\right|\leq\tau\right\}\widehat{s}_{N,t}^{\alpha}(\theta)\big{(}\widehat{s}_{N,s}^{\alpha}(\theta)\big{)}^{\prime},$
$\widehat{s}_{i,t}^{\alpha}(\theta)$ $:=\bigg{[}\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)-\frac{1}{T}\sum_{t=1}^{T}\partial_{\alpha_{i}}l_{i,t}(\theta,\phi)\bigg{]}\bigg{|}_{\phi=\widehat{\phi}(\theta)}.$
Here, $\tau$ is a truncation parameter commonly used in the literature (see, e.g., Hahn and Kuersteiner (2011)). When the data exhibit serial dependence, we set $\tau=1$ or $\tau=2$ as in Fernández-Val and Weidner (2016). Our simulations show that the results for $\tau=1$ and $\tau=2$ are similar. For data without serial dependence, we set $\tau=0.$ Additionally, we define
$\widehat{\Sigma}^{\breve{\gamma}}_{\nu}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathds{1}\left\{\left|i-j\right|\leq\upsilon\right\}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{(}\widehat{S}^{\breve{\gamma}}_{j}(\theta)\big{)}^{\prime}$
$\widehat{S}^{\breve{\gamma}}_{i}(\theta)$ $:=\big{(}(\widehat{s}_{i,1}^{\gamma}(\theta))^{\prime},\ldots,(\widehat{s}_{i,T-1}^{\gamma}(\theta))^{\prime}\big{)}^{\prime}-\iota_{T-1}\otimes\widehat{s}_{i,T}^{\gamma}(\theta)$
$\widehat{s}_{i,t}^{\gamma}(\theta)$ $:=\bigg{[}\partial_{\gamma_{t}}l_{i,t}\left(\theta,\phi\right)-\frac{1}{N}\sum_{i=1}^{N}\partial_{\gamma_{t}}l_{i,t}\left(\theta,\phi\right)\bigg{]}\bigg{|}_{\phi=\widehat{\phi}(\theta)},\quad i=1,\ldots,N,\ t=1,\ldots,T.$
Specifically, when we set $\upsilon=0$ given conditional cross-sectional independence, $\widehat{\Sigma}^{\breve{\gamma}}_{\nu}(\theta)$ simplifies to
@@
$\widehat{\Sigma}^{\breve{\gamma}}_{0}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{(}\widehat{S}^{\breve{\gamma}}_{i}(\theta)\big{)}^{\prime}.$
Thus, using the estimated bias, we obtain the corrected likelihood $L(\theta):$
$L(\theta)=\widehat{l}(\theta)+\widehat{B}_{\breve{\alpha}}(\theta)+\widehat{B}_{\breve{\gamma}}(\theta)$
as defined in (2.11).
The following theorem demonstrates that the difference between the corrected likelihood $L(\theta)$ and and the infeasible likelihood $l(\theta)$ is of order $o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$ implying that $L(\theta)$ no longer suffers from IPP.
@@
Under Assumption 1, the corrected likelihood satisfies
$L(\theta)=l(\theta)+o_{\mathbb{P}}\big{(}1/\sqrt{NT}\big{)},$
as $N$ and $T\to\infty.$
Compared to $\widehat{l}(\theta),$ the difference between the corrected likelihood and the infeasible likelihood is negligible. The corrected likelihood is derived from the expansion (2.7). Specifically, (2.7) is based on two Taylor expansions: the expansion of the score function for incidental parameters in (2.3) and the expansion of the original likelihood in (2.6). It is seen that the bias terms originate from the quadratic form on the right-hand side of (2.7), expressed in terms of the score and Hessian of the incidental parameters. Under Assumption 1, the quadratic form can be further decomposed into two additive sub-quadratic forms (Lemma 1), capturing the bias contributions from the individual and time dimensions, respectively. Consequently, our bias-corrected likelihood represents the original likelihood augmented by the estimation of these two additive bias terms.
Last, we demonstrate that the MLE obtained from the corrected likelihood, $\widehat{\theta}_{L}=\arg\max_{\theta}L(\theta),$ follows an limiting normal distribution with zero mean. Additional regularity conditions are required.

### v4

@@ 3 Bias Correction and Asymptotic Theory
$s(\theta)-\mathbb{E}s\left(\theta\right)=:\left(\begin{array}[c]{c}\widetilde{s}_{\alpha}(\theta)\\ \widetilde{s}_{\gamma}(\theta)\end{array}\right),\qquad[\mathbb{E}H(\theta)]^{-1}=:\left(\begin{array}[c]{cc}H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&H_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right),$ (3.2)
where $\widetilde{s}_{\alpha}(\theta)$ is $(N-1)K\times 1$ and $\widetilde{s}_{\gamma}(\theta)$ is $(T-1)K\times 1$ while $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ is $(N-1)K\times(N-1)K$ and $H_{\gamma\gamma}^{\ast}(\theta)$ is $(T-1)K\times(T-1)K$. The off-diagonal blocks $H_{\alpha\gamma}^{\ast}(\theta)$ and $H_{\gamma\alpha}^{\mathcal{\ast}}(\theta)$ are defined accordingly. Further denote
$S_{\alpha\alpha}(\theta):=\mathbb{E\{}\widetilde{s}_{\alpha}(\theta)\mathbb{[}\widetilde{s}_{\alpha}(\theta)]^{\prime}\},\qquad S_{\gamma\gamma}(\theta):=\mathbb{E\{}\widetilde{s}_{\gamma}(\theta)\mathbb{[}\widetilde{s}_{\gamma}(\theta)]^{\prime}\}.$
The next theorem formally states the asymptotic expansion of $\widehat{l}(\theta)$.
@@
Under Assumption 1 and uniformly over $\theta\in\Theta$,
$\mathbb{E}l(\theta)=\mathbb{E}\widehat{l}(\theta)+B_{\alpha}(\theta)+B_{\gamma}(\theta)+o(1/\sqrt{NT}),$
as $N,T\rightarrow\infty$, where
@@
$B_{\alpha}(\theta):=\frac{1}{2}\operatorname*{trace}[S_{\alpha\alpha}(\theta)H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad B_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}S_{\gamma\gamma}(\theta)H_{\gamma\gamma}^{\ast}(\theta)],$
and $B_{\alpha}(\theta)+B_{\gamma}(\theta)=O(1/\sqrt{NT})$.
[…]
The bias terms $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ involves $\psi(\theta)$ and $\mathbb{E}$ which are infeasible to compute. However, $B_{\alpha}(\theta)$ and $B_{\gamma}(\theta)$ can be estimated by plug-in estimates of $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, $H_{\gamma\gamma}^{\ast}(\theta)$, $S_{\alpha\alpha}(\theta)$, and $S_{\gamma\gamma}(\theta)$. The two Hessian terms can be estimated in a straightforward way. Denote $\widehat{H}(\theta):=H(\theta,\widehat{\psi}(\theta))$ and partition $[\widehat{H}(\theta)]^{-1}$ into
@@
$[\widehat{H}(\theta)]^{-1}=:\left(\begin{array}[c]{cc}\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\alpha\gamma}^{\mathcal{\ast}}(\theta)\\ \widehat{H}_{\gamma\alpha}^{\mathcal{\ast}}(\theta)&\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)\end{array}\right)$ (3.3)
in the same manner as in (3.2). The term $H_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$ can be estimated by the $(N-1)K\times(N-1)K$ block $\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)$, and $H_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$ by the $(T-1)K\times(T-1)K$ block $\widehat{H}_{\gamma\gamma}^{\mathcal{\ast}}(\theta)$. Estimators of $S_{\alpha\alpha}(\theta)$ and $S_{\gamma\gamma}(\theta)$ are constructed from the derivatives of $l_{i,t}(\theta,\phi)$, the original likelihood for individual $i$ at time $t$. This procedure needs some explanation. The reparameterization ensures the following identity:
$S_{\alpha\alpha}(\theta)=D_{1}\mathcal{S}_{\alpha\alpha}(\theta)D_{1}^{\prime},$
where $\mathcal{S}_{\alpha\alpha}(\theta)$ is the outer product of the $NK\times 1$ vector of $\partial_{\alpha}\ell(\theta,\phi)-\mathbb{E}\partial_{\alpha}\ell(\theta,\phi)$ evaluated at $\phi=\phi(\theta)=D^{\prime}\psi(\theta)$, where $\partial_{\alpha}\ell(\theta,\phi)$ is the derivative of the original likelihood $\ell(\theta,\phi)$ w.r.t. $\alpha:=(\alpha_{1}^{\prime},\ldots,\alpha_{N}^{\prime})^{\prime}$. Specifically,
$\mathcal{S}_{\alpha\alpha}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E\{}\widetilde{s}_{t}^{\alpha}(\theta)[\widetilde{s}_{s}^{\alpha}(\theta)]^{\prime}\},\qquad\widetilde{s}_{t}^{\alpha}(\theta):=\{[\widetilde{s}_{1,t}^{\alpha}(\theta)]^{\prime},\ldots,[\widetilde{s}_{N,t}^{\alpha}(\theta)]^{\prime}\}^{\prime},$
$\widetilde{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\alpha}(\theta,\phi(\theta)),\qquad s_{i,t}^{\alpha}(\theta,\phi):=\partial_{\alpha_{i}}l_{i,t}(\theta,\phi).$
Here, $\widetilde{s}_{i,t}^{\alpha}(\theta)$ is the centered version of $s_{i,t}^{\alpha}(\theta,\phi(\theta))$, the derivative, w.r.t. $\alpha_{i}$ and evaluated at $\phi(\theta)$, of the original likelihood $l_{i,t}(\theta,\phi)$ for individual $i$ at time $t$. Under cross-sectional independence, $\mathcal{S}_{\alpha\alpha}(\theta)$ simplifies to
@@
$\mathcal{S}_{\alpha\alpha}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathbb{E}\operatorname*{diag}\{\widetilde{s}_{1,t}^{\alpha}(\theta)[\widetilde{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widetilde{s}_{N,t}^{\alpha}(\theta)[\widetilde{s}_{N,s}^{\alpha}(\theta)]^{\prime}\}.$
Naturally, the estimator of $S_{\alpha\alpha}(\theta)$ is
$\widehat{S}_{\alpha\alpha}(\theta)$ $:=D_{1}\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)D_{1}^{\prime},$ (3.4)
@@
$\widehat{\mathcal{S}}_{\alpha\alpha}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{t=1}^{T}\sum_{s=1}^{T}\mathds{1}\{\left|t-s\right|\leq\tau\}\operatorname*{diag}\{\widehat{s}_{1,t}^{\alpha}(\theta)[\widehat{s}_{1,s}^{\alpha}(\theta)]^{\prime},\ldots,\widehat{s}_{N,t}^{\alpha}(\theta)[\widehat{s}_{N,s}^{\alpha}(\theta)]^{\prime}\},$
$\widehat{s}_{i,t}^{\alpha}(\theta)$ $:=s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\alpha}(\theta,\widehat{\phi}(\theta)).$
The indicator $\mathds{1}\{\left|t-s\right|\leq\tau\}$ is a truncation mechanism with truncation parameter $\tau$ (see Remark 3 below). Similarly for $S_{\gamma\gamma}(\theta)$, the reparameterization produces the identity:
$S_{\gamma\gamma}(\theta)=D_{2}\mathcal{S}_{\gamma\gamma}(\theta)D_{2}^{\prime},$
where $\mathcal{S}_{\gamma\gamma}(\theta)$ is the outer product of the $TK\times 1$ vector of $\partial_{\gamma}\ell(\theta,\phi)-\mathbb{E}\partial_{\gamma}\ell(\theta,\phi)$, for $\gamma:=(\gamma_{1}^{\prime},\ldots,\gamma_{T}^{\prime})^{\prime}$, evaluated at $\phi=\phi(\theta)$:
$\mathcal{S}_{\gamma\gamma}(\theta)$ $=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\sum_{j=1}^{N}\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\},\qquad\widetilde{s}_{i}^{\gamma}(\theta):=\{[\widetilde{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widetilde{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
$\widetilde{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\phi(\theta))-\mathbb{E}s_{i,t}^{\gamma}(\theta,\phi(\theta)),\qquad s_{i,t}^{\gamma}(\theta,\phi):=\partial_{\gamma_{t}}l_{i,t}(\theta,\phi),$
where $\widetilde{s}_{i,t}^{\gamma}(\theta)$ is the centered version of $s_{i,t}^{\gamma}(\theta,\phi(\theta))$, the derivative of $l_{i,t}(\theta,\phi)$ w.r.t. $\gamma_{t}$ and evaluated at $\phi(\theta)$. By cross-sectional independence, $\mathcal{S}_{\gamma\gamma}(\theta)$ similarly reduces to
$\mathcal{S}_{\gamma\gamma}(\theta)=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}s_{i}^{\gamma}(\theta)[s_{i}^{\gamma}(\theta)]^{\prime}\}.$ (3.5)
Intuitively, the estimator for $S_{\gamma\gamma}(\theta)$ is
$\widehat{S}_{\gamma\gamma}(\theta)$ $=D_{2}\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)D_{2}^{\prime},$ (3.6)
@@
$\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$ $:=\frac{1}{N^{2}T^{2}}\sum_{i=1}^{N}\mathbb{E\{}\widehat{s}_{i}^{\gamma}(\theta)[\widehat{s}_{i}^{\gamma}(\theta)]^{\prime}\},\qquad\widehat{s}_{i}^{\gamma}(\theta):=\{[\widehat{s}_{i,1}^{\gamma}(\theta)]^{\prime},\ldots,[\widehat{s}_{i,T}^{\gamma}(\theta)]^{\prime}\}^{\prime},$
$\widehat{s}_{i,t}^{\gamma}(\theta)$ $:=s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta))-\frac{1}{T}\sum_{t=1}^{T}s_{i,t}^{\gamma}(\theta,\widehat{\phi}(\theta)).$
Remark 3 (truncation)
In the estimator $\widehat{S}_{\alpha\alpha}(\theta)$, we use the truncation technique commonly found in e.g., Hahn and Kuersteiner (2011) (see also Hahn and Kuersteiner (2007)), Fernández-Val and Weidner (2016), and Galvao and Kato (2016). This is related to the literature on consistent estimations of autocovariances, $\mathbb{E\{}\widetilde{s}_{i,t}^{\alpha}(\theta)[\widetilde{s}_{i,s}^{\alpha}(\theta)]\}$ in our case. Detailed discussions in this regard can be found in Andrews (1991). Without truncation, we would have $\widehat{S}_{\alpha\alpha}(\theta)=\widehat{s}_{\alpha}(\theta)[\widehat{s}_{\alpha}(\theta)]^{\prime}=0$ for every $\theta$ because $\widehat{s}_{\alpha}(\theta)=0$ by definition, where $\widehat{s}_{\alpha}(\theta)$ is defined from $\widehat{s}(\theta):=s(\theta,\widehat{\psi}(\theta))$ under the same partition in (3.2). Here, the truncation parameter $\tau\rightarrow\infty$ but sufficiently slower relative to $T$ (see Theorem 2 below). Hahn and Kuersteiner (2011) suggests the default value of $\tau=1$ for small $T$. For intermediate $T$, we recommend following the advices from Fernández-Val and Weidner (2016) to conduct a sensitivity test for different values of $\tau$ and to avoid using very large $\tau$ values. In our simulation, we use $\tau=1$ and $2$ and find the difference relatively insignificant.
On the other hand, the term $\widehat{S}_{\gamma\gamma}(\theta)$ does not require any truncation, because we have already eliminated $\mathbb{E\{}\widetilde{s}_{i}^{\gamma}(\theta)[\widetilde{s}_{j}^{\gamma}(\theta)]^{\prime}\}$ for $i\neq j$ from Equation (3.5) using cross-sectional independence. One may still opt to truncate $\widehat{\mathcal{S}}_{\gamma\gamma}(\theta)$, such that all elements beyond the $\tau$-th sub- and super-diagonal are symmetrically set to 00. We have experimented with such truncations but found little impact on $\widehat{\theta}_{L}$ and test statistics. The comparison is, hence, not reported in the paper.
Using Equations (3.3), (3.4), and (3.6), we can construct the corrected likelihood $L(\theta)$, which is stated in the theorem below with its rate of convergence.
@@
Under Assumption 1 and for some $\tau\rightarrow\infty$ such that $\tau/T\rightarrow 0$,
$L(\theta)-l(\theta)=o_{\mathbb{P}}(1/\sqrt{NT}),$
uniformly over $\theta\in\Theta$ as $N,T\rightarrow\infty$, where
$L(\theta):=\widehat{l}(\theta)+\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta),$
as in (3.1), with
$\widehat{B}_{\alpha}(\theta):=\frac{1}{2}\operatorname*{trace}[\widehat{S}_{\alpha\alpha}(\theta)\widehat{H}_{\alpha\alpha}^{\mathcal{\ast}}(\theta)],\qquad\widehat{B}_{\gamma}(\theta):=\frac{1}{2}\operatorname*{trace}\mathbb{[}\widehat{S}_{\gamma\gamma}(\theta)\widehat{H}_{\gamma\gamma}^{\ast}(\theta)].$
As mentioned above, Theorem 2 states that $L(\theta)$ does not suffer from the IPP. Showing 2 essentially amounts to establishing $\widehat{B}_{\alpha}(\theta)+\widehat{B}_{\gamma}(\theta)=B_{\alpha}(\theta)+B_{\gamma}(\theta)+o_{\mathbb{P}}(1/\sqrt{NT})$ under $\tau/T\rightarrow 0$ and Assumption 1. Here, the difference between $\widehat{B}_{\alpha}(\theta)$ and $B_{\alpha}(\theta)$ is due to the truncation and the estimation error. For the truncation, we show that $1/\left(N^{2}T^{2}\right)\sum_{t=1}^{T}\sum_{s=1}^{T}\left(\mathds{1}\{\left|t-s\right|>\tau\}\cdot\right)$, which is the truncated (i.e., the removed) part in (3.4), is negligible relative to $1/\sqrt{NT}$ under the aforementioned conditions. Subsequently, the uniform law of large numbers guarantees that the estimation error is also negligible relative to $1/\sqrt{NT}$.
Next, we turn our attention to the asymptotic properties of the $L(\theta)$-based estimators $\widehat{\theta}_{L}=\arg\max_{\theta}L(\theta)$ and of the $L(\theta)$-based LR, LM, Wald test statistics. We enforce the following assumption.